Find the values of the variables in DEFG
DG= 2x+2
EF= 3x-3
DE= 3x+1
GF= 2x+6

Answer by ankor@dixie-net.com(22740)

Find the values of the variables in DEFG
DG= 2x+2
EF= 3x-3
DE= 3x+1
GF= 2x+6
Assuming opposite sides are equal
EF = DG
therefore
3x-3 = 2x+2
3x - 2x = 2 + 3
x = 5
:
You can check this using DE = GF